factorize:27x^3+8x^3+1

Grade:10

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Its (3x)^3+(2x)^3+1^3

use the expansion for a3+b3+c3

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sudheesh Singanamalla
13 years ago

27 x^3 + 8 x^3 + 1

= 35 x^3 + 1

a3 + b3  = (a + b)(a2 – ab + b2)

so

35 x^3 + 1 = (35x + 1 ) ( 35 x^2 - 35 x + 1)

 Solving the (35x^2 - 35x + 1) as quadratic equation)

 roots are 35 + sqrt(1085) and 35 - sqrt(1085)

35x+1 = 0 ----------

==> x = -1/2

therefore roots are

x = -1/2

x = 35 + sqrt (1085)

x = 35 - sqrt (1085)

--------------------------------------------------------------------------------------

PLEASE APPROVE

--------------------------------------------------------------------------------------

indrayani bargaje
13 years ago

pleaSE tell me the identity of a^3+b^3+c^3